OpenStudy (anonymous):

If a≠b and (1/x) + (1/a) = (1/b), then x= How does "x" end up being (ab/a-b)?

OpenStudy (asnaseer):

subtract 1/a from both sides - what do you end up with?

OpenStudy (anonymous):

If a≠b and (1/x) + (1/a) = (1/b), then (1/x) = (1/b) - (1/a), right?

OpenStudy (anonymous):

Right

OpenStudy (asnaseer):

now simplify 1/b - 1/a

OpenStudy (anonymous):

so how do simplify the RHS of the expression?

OpenStudy (anonymous):

RHS=right hand side

OpenStudy (anonymous):

Factor out the - ... or create a common denominator?

OpenStudy (anonymous):

right,create common denominator, so whats the LCM

OpenStudy (anonymous):

(a+b)/ab

OpenStudy (anonymous):

LCM is obtained by multiplying the two denominators, i.e a*b = ab, now with this can you solve (1/b) - (1/a)

OpenStudy (anonymous):

Made a mistake by factoring out the - in my work. I have worked out the problem and got the correct answer. Thank you for the help.

OpenStudy (anonymous):

good